A steady flow, steady state thermodynamic system receives 45 kg/min of a fluid at 207kPa and discharges it from point 24 m. above the entrance section at 1034 kPa. The fluid enters with a velocity of 36.6 m/s and leaves with a velocity of 12 m/s. During this process, there are supplied 7 kW of heat from an external source, and the increase in enthalpy is 4.6 kJ/kg.


1
Expert's answer
2021-10-31T18:16:28-0400

"wf_1=p_1V_1=5.351~\frac{kJ}{kg},"

"wf_2=p_2V_2=106.459~\frac{kJ}{kg},"

"KE_1=\frac{v_1^2}{2k}=7.442~\frac{kJ}{kg},"

"KE_2=\frac{v_2^2}{2k}=16.745~\frac{kJ}{kg},"

"E_{in}=E_{out},"

"U_1+wf_1+KE_1+Q=U_2+wf_2+KE_2+W,"

"W=-486~\frac{kJ}{min}."